OpenStudy (anonymous):

Find a ordered pair for... 3x-y=3

OpenStudy (anonymous):

@sudonick @csscgirl

OpenStudy (anonymous):

three more :D (0,-3),(1,0),(2,3)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

lol i may b n love with that sudonick person :P XD

OpenStudy (anonymous):

x-2y = 8 ?

OpenStudy (anonymous):

lol @csscgirl

OpenStudy (anonymous):

(10,1)

OpenStudy (anonymous):

(2,-3) andddd

OpenStudy (anonymous):

(8,0)

OpenStudy (anonymous):

(0,-4)

OpenStudy (anonymous):

snapp

OpenStudy (anonymous):

last one.. 6x-y=9

OpenStudy (anonymous):

(0,-9),(1,-3),(2,3)

OpenStudy (anonymous):

(3,9)

OpenStudy (phi):

if you want to find a point that works for x-2y = 8 replace y with 0 (because -2*0 is 0, and you get x = 8 the answer is (8,0)

OpenStudy (anonymous):

Thanks @csscgirl @phi @sudonick

OpenStudy (anonymous):

yvw :)